OpenStudy (becks):

A six-member committee is formed to plan a Christmas formal. The members of the group are to be selected from five teachers and nine students. If the members are randomly selected, what is the probability that it will include at least two teachers? How do I figure out the probability!?!? :O) HELP, PLZ!

OpenStudy (mathmate):

Total number of different committees = C(14,6) = N We have two cases where we have less than 2 teachers. case 1: 1 teacher (out of 5) Number of ways to choose 1 teacher = C(5,1) = 5 Number of ways to choose the remaining members from students = C(9,5) Total number of ways to choose a committee with exactly 1 teacher = 5C(9,5) = N1 case 2: no teacher Number of ways to choose 6 students = C(9,6) = N2 Probability to choose at least two teachers = (N-N1-N2)/N =109/143

OpenStudy (becks):

^ Thanks, SOOO much! This was a major help. :DD !!

OpenStudy (mathmate):

You're welcome! :)